An elevator starts from rest with a constant upward acceleration.It moves 2m in the forst 0.6 sec.A passanger in the elevator is holding a 3kg package by a vertical string.What is the tension in the string during acceleration.(take g=9.8 m/s^2)


Grade:10

10 Answers

B.V.Suhas Sudheendhra
12 years ago

Here find the acceleration of the elevator by using the formula : s=ut + 0.5at2 

where u is zero(as it is rest initially) t = 0.6s, s=2m.

then draw the free body diagram for the block.

the weight is acting downwards=mg (initially when the lift is at rest)

since the acceleration of lift is in the upward direction = a

therefore the relative acceleration is a1=(g+a)

then tension in the string T = ma1

Substitute the values to get the answer.

 

Swapnil Saxena
12 years ago

The elevator in this problem is an non inertial frame, and so newtons laws can be directly applied to it.

Pseudo forces should also be taken into consideration.

For the pseudo forces, the acceleration of the block can be found by the following method

= 2 = 1/2(a)(0.6)2

= 4=(0.36)a

=a= (4/0.36)=11.11

Then pseudo forec = 3*(4/0.36)=33.33

thn the tension in the string = 30N(weight of the body)+33.33N(pseudo force)

= 63.33N

anand .
12 years ago

calculate acceleration s=1/2at2

s=2

t=0.6

a=11.1

then drawing fbd

t=ma+mg

t=62.7newton

Sathya
12 years ago

@ Swapnil saxena,What is pseudo force?

To all,

I have a doubt,

the acceleration will be acting upwards , the gravitational acceleration and tension will be acting downwards .

Thus it should be mg+t=ma ,na?

How is mg+ma=t ???

Ashwin Muralidharan IIT Madras
12 years ago

Hi Satyaram,

 

Net Force = Mass*Acceleration.

Hence T - mg = ma.

Net force in the upward Direction = T - mg, which should be mass*acn [Newtons Law].

Please dont question Newtons Law.

Assume it is valid.

 

Regards,

Ashwin (IIT MadraS).

Sathya
12 years ago

But,always,the tension will be in the direction opp. to direction of  motion of object,right?

Ashwin Muralidharan IIT Madras
12 years ago

Hi Sathyaram,

 

Is that so... Just think about it.

 

Say we have a pulley, with two masses hanging through a rope on either sides of the pulley. Now the heavier mass will move down and the lighter mass will move up.

But still the tension in the rope for both the masses will be in the upward direction.

So it is not required that the tension will be in the opposite direction of motion of the object.

 

Regards,

Ashwin (IIT Madras).

Sathya
12 years ago

Sorry,I take back my words....

What is pseudo force?

Ashutosh
6 years ago
Pseudo force is a force that applies on a body which we cannot see but it does come into effect this so called pseudo force Is explained by the common example when a person moves in a merry go round the rope clings to the other side this.is called centrifugal force a pseudo forceSimilarly u cannot see or prove normal reaction but we must assume that there is a force that acts or when a box is kept on a table it wud simply break which it does notHope u.did.understand Regard Class 9 ashutosh
kanishka
6 years ago
Pseudo force is a force that applies on a body which we cannot see but it does come into effect this so called pseudo force Is explained by the common example when a person moves in a merry go round the rope clings to the other side this.is called centrifugal force a pseudo forceSimilarly u cannot see or prove normal reaction(weight) but we must assume that there is a force that acts or when a box is kept on a table it wud simply break.